. A battery of e.m.f 6.0 V passes a current of 0.3 A through a torch bulb for 5 minutes. How much energy is transferred from the cell?
___________________________
1
Expert's answer
2020-06-25T09:51:37-0400

Electric energy can be calculated according to:

"A=I*U*t"

For our case I = 0.3 A, U = 6 V, t = 5 minutes = 5*60 = 300 seconds.


So we have:

"A=0.3A*6V*300sec=540J"


Answer: 540J of energy was transferred from the cell.